# Independent sets of non-geometric lattices and the maximal adjoint

### v2

@@ Abstract.
We present a construction of a family of independent sets for a finite, atomic, and graded lattice generalizing the known cryptomorphism between geometric lattices and matroids. This family then gives rise to an embedding theorem into geometric lattices that preserve the set of atoms. We then apply these theorems to the concept of adjoint matroids, obtaining a characterization and proving a conjecture on the combinatorial derived matroid for uniform and co-rank 3 matroids. Lastly, we use our characterization of adjoints to compute the list of adjoints of some simple examples.
@@ 1. Introduction
In [10] Andreas Blass and Bruce E. Sagan introduced the concept of $NBB$ sets as a way to calculate the möbius function of a finite and bounded lattice by ordering the atoms of the lattice. These sets can be thought of as a generalization of $NBC$ sets for ordinary matroids. In [9] the family of $NBB$ sets on all linear orders on the atoms of the lattice was used to form a family of independent sets and show that if it were a family of sets of a matroid, then the lattice could be order-embedded into the corresponding geometric lattice. We extend these results here and apply them to the construction of the adjoint matroid.
The structure of this paper is as follows. Section 2 is dedicated to preliminaries on lattices,matroids, and the adjoint matroid. In section 3 we discuss the family of independent sets of a finite, atomic and graded lattice, recalling results from [9] and presenting a new embedding theorem. In section 4 we show the effects of basic lattice operations on its family of independent sets and a class of special maximal independent sets. Section 5 moves to working with adjoint matroids centered on the characterization theorems of adjoints as an application of the family of independent sets. In Section 6 we present some partial results showing that the combinatorial derived matroid is an adjoint and the existence of a maximal adjoint. Finally, this paper also includes an appendix in which we formulate an algorithm for checking for adjoints based on the theorems from Section 5. We then use this algorithm to calculate the list of adjoints of some matroids of small rank and co-rank.
@@ 2. Preliminaries
We start with some basic lattice properties; for a more extensive background on matroids and geometric lattices, see [1]. A partially ordered set $\mathcal{L}$ in which every pair of elements has a unique supremum (also called join and denoted $\vee$) and a unique infimum (also called meet and denoted $\wedge$) is called a lattice. We say that a lattice is bounded if it has a greatest element (denoted $\hat{1}$) and a least element (denoted $\hat{0}$).
@@ 2.1. Definition:
We say $B$ covers $A$ in $\mathcal{L}$ and denote $A\lessdot B$ if $A<B$ and for all $C$ such that $A\leq C\leq B$ we have $C=B$ or $C=A$. $A$ is an atom if it covers $\hat{0}$ and a coatom if it is covered by $\hat{1}$. A bounded lattice is called atomic if every element is a join of atoms and coatomic if every element is a meet of coatoms. We denote the set of atoms by $A(\mathcal{L})$.
Finally, we say that a bounded lattice is graded if there exists a rank function $rank:\mathcal{L}\rightarrow\mathbb{N}$ with the following properties:
@@
    Compatibility with the order of $\mathcal{L}$, $A\leq B\iff rank(A)\leq rank(B)$.
@@
    Compatibility with the covering relation of $\mathcal{L}$, $A\lessdot B\Rightarrow rank(A)=rank(B)+1$.
@@ 2.2. Definition:
A geometric lattice is a finite graded lattice such that $rank(A)+rank(B)\leq rank(A\vee B)+rank(A\land B)$.
Moving on to some basic matroid properties (for more details, see [1]).
We will follow the standard conventions, a matroid $\mathcal{M}$ will refer to the set of bases of the matroid, and we denote by $I(\mathcal{M}),\mathcal{C}(\mathcal{M}),\mathcal{L}(\mathcal{M}),E(\mathcal{M})$ the family of independent sets, the collection of circuits, the corresponding geometric lattice, and the ground set accordingly. When it is obvious which matroid we refer to, we will omit $\mathcal{M}$, for example $\mathcal{L}$ instead of $\mathcal{L}(\mathcal{M})$.
An element $a\in E$ will be called a loop of $\mathcal{M}$ if $i\notin I$ for every $I\in\mathcal{M}$. Two elements $a,b\in E$ will be called parallel if both are not loops and $\left\{a,b\right\}\notin I$ for every $I\in\mathcal{M}$. We will say that a matroid is simple if it has no loops or parallel elements.
@@ 2.3. Definition:
Given two matroids $\mathcal{M}_{1},\mathcal{M}_{2}$ on the ground set $E$, we say $\mathcal{M}_{1}\leq\mathcal{M}_{2}$ in the weak order of matroids if $\mathcal{M}_{1}\subseteq\mathcal{M}_{2}$.
The dual matroid of $\mathcal{M}$ denoted $\mathcal{M}^{*}$ is the matroid on the same ground set, and in which a set is independent if and only if $\mathcal{M}$ has a basis set disjoint from it. We denote the geometric lattice corresponding to $\mathcal{M}^{*}$ by $\mathcal{L}^{*}$ and a loop of $\mathcal{M}^{*}$ will be called a co-loop.
A known property is that the complements of the hyperplanes (coatoms of the geometric lattice) are precisely the circuits of the dual matroid. This property allows us to view an extension of a dual lattice of a geometric lattice discussed in the following sections as a matroid with ground set $\mathcal{C}(\mathcal{M})$. We formulate our definitions with this change in mind:
@@ 2.4. Definition:
Let $\mathcal{L}$ be a geometric lattice and $\mathcal{L}^{opp}$ its dual poset. We say $\mathcal{L}^{\triangle}$ is an adjoint of $\mathcal{L}^{*}$ if there exists an extension of a bijection on the set of atoms $f^{\prime}:A(\mathcal{L}^{opp})\rightarrow A(\mathcal{L}^{\triangle})$ to a rank preserving order embedding $f:\mathcal{L}^{opp}\rightarrow\mathcal{L}^{\triangle}$. We say $\mathcal{M}^{\triangle}$ is an adjoint matroid of $\mathcal{M}$ if the corresponding geometric lattice of $\mathcal{M}^{\triangle}$ is an adjoint of $\mathcal{L}^{*}$.
The following two properties are known for adjoints (for example, in [5])
@@ 3.1. Definition:
Let $(\mathcal{L},\omega)$ be a pair of a finite bounded lattice $\mathcal{L}$ and a partial order $\omega$ on its set of atoms $A(\mathcal{L})$. A nonempty set $D\subseteq A(\mathcal{L})$ of atoms is bounded below or $BB$ if there exist $a\in A(\mathcal{L})$ such that $a$ is a strict lower bound for all $d\in D$ in the order $\omega$ and $a\leq\vee D$ in $\mathcal{L}$.
@@
Given a bounded lattice $\mathcal{L}$ and $P$ the set of all linear orders on $A(\mathcal{L})$ we define the family of independent sets $I(\mathcal{L})$ of $\mathcal{L}$ as follows:
$I\left(\mathcal{L}\right)=\left\{A\in P\left(A(\mathcal{L})\right)\mid\exists\omega\in P\text{ s.t. }A\text{ is an }NBB\text{ set in }\left(\mathcal{L},\omega\right)\right\}$
We refer to elements in $I(\mathcal{L})$ as independent sets of $\mathcal{L}$; note that sets containing at most two atoms are always independent. The constructions in the section can be thought of as a generalization of the following observation.
@@ 3.2. Theorem:
If $\mathcal{L}$ is a geometric lattice then $I(\mathcal{L})$ is the family of independent sets of the matroid corresponding to $\mathcal{L}$.
For the remainder of this section, we fix a finite, bounded, atomic, and graded lattice $\mathcal{L}$. If $\vee A=X\in\mathcal{L}$ for a set of atoms $A\subseteq A(\mathcal{L})$ we say that $A$ spans $X$ and $rank(A)\coloneqq rank(X)$.
@@ 3.3. Definition:
An independent set $I$ of $\mathcal{L}$ is called geometric if $rank(I)=\left|I\right|$.
It was shown in [9] that the size of an independent set is always $\leq$ then its rank and that every non-geometric independent set can be extended to one of the same rank, consequently the rank of an element equals the size of a maximal independent set contained in it. The first embedding theorem was introduced in [9] and covers the case in which $I(\mathcal{L})$ is a family of independent sets of a matroid.
@@ 3.4. Theorem:
If $I(\mathcal{L})$ is a family of independent sets of a matroid, then there exists a rank-preserving order embedding $f:\mathcal{L}\rightarrow\mathcal{P}$ for which the restriction to the set of atoms $f\mid_{A(\mathcal{L})}:A(\mathcal{L})\rightarrow A(\mathcal{P})$ is a bijection. With $\mathcal{P}$ being the geometric lattice corresponding to $I(\mathcal{L})$.
We continue with a formulation of an embedding theorem in the case where $I(\mathcal{L})$ is not a family of independent sets of a matroid.
@@ 3.5. Lemma:
Let $f^{\prime}:A(\mathcal{L})\rightarrow A(\mathcal{P})$ be a bijection between the set of atoms of $\mathcal{L}$ and a geometric lattice $\mathcal{P}$. If $f^{\prime}$ can be extended to a rank-preserving order embedding $f:\mathcal{L}\rightarrow\mathcal{P}$, then $f(I(\mathcal{L}))\subseteq I(\mathcal{P})$.
@@
Let $f:\mathcal{L}\rightarrow\mathcal{P}$ such an embedding and choose a minimal $I\in f(I(\mathcal{L}))\setminus I(\mathcal{P})$. Let $a\in f^{-1}(I)$ with $rank(f^{-1}(I)\setminus\left\{a\right\})\lneq rank(f^{-1}(I))$ guaranteed by the definition of $NBB$ sets. As $I\setminus\left\{f(a)\right\}$ is an independent set of $\mathcal{P}$, we must have
$f(a)\in\vee_{\mathcal{P}}(I\setminus\left\{f(a)\right\})\leq_{\mathcal{P}}f(\vee_{\mathcal{L}}(I\setminus\left\{f(a)\right\}))$
which is a contradiction to the choice of $a$.
3.6. Corollary:
Let $\mathcal{L}$ be a finite graded lattice (not necessarily atomic) and $\mathcal{P}$ be a geometric lattice. If there exists a rank-preserving order embedding $f:\mathcal{L}\rightarrow\mathcal{P}$ with $f\mid_{A(\mathcal{L})}$ an injection into $A(\mathcal{P})$, then $f(I(\mathcal{L}))\subseteq I(\mathcal{P})$.
@@ 3.7. Theorem:
Let $f^{\prime}:A(\mathcal{L})\rightarrow A(\mathcal{P})$ a bijection between the set of atoms of $\mathcal{L}$ and the set of atoms of a geometric lattice $\mathcal{P}$ of the same rank. $f^{\prime}$ can be extended to a rank-preserving order embedding $f:\mathcal{L}\rightarrow\mathcal{P}$ iff $f(I(\mathcal{L}))\subseteq I(\mathcal{P})$ and $f(A)\notin I(\mathcal{P})$ for every $A\in 2^{A(\mathcal{L})}$ such that $\left|A\right|\lneq rank\left(\vee A\right)$.
@@
If there exists a function $f$ then Lemma [3.5] implies the first condition. The second condition holds as if $f(A)\in I(\mathcal{P})$ then $rank(f(A))=\left|A\right|$ and $f$ is rank-preserving.
In the other direction, we define $f(X)\coloneqq\vee_{\mathcal{P}}(f(A(X)))$. First, notice that $rank(X)\leq rank(f(X))$ for all $X\in\mathcal{L}$ as $f(X)$ contains the image of a geometric spanning set of $X$. To see $rank(X)\geq rank(f(X))$ for all $X\in\mathcal{L}$ observe that it cannot contain a larger independent set of $\mathcal{P}$ as it will contradict the second condition. To see $f$ is an embedding, let $X\in\mathcal{L}$, $a\in A(\mathcal{L})\setminus A(X)$ and $B$ be a geometric independent set spanning $X$. Therefore, $f(A)\cup\left\{f(a)\right\}\in I(\mathcal{P})$ is an independent set of size $rank(X)+1$ in $f(X)$ which contradicts $f$ being rank-preserving.

### v3

@@ Abstract.
We construct a family of independent sets for finite, atomic, and graded lattices, extending the well-known cryptomorphism between geometric lattices and matroids. This construction leads to an embedding theorem into geometric lattices that preserves the set of atoms. We then apply these results to adjoint matroids, providing new characterizations of adjoints and partially proving a conjecture on the combinatorial derived matroid. Finally, we use our characterization of adjoints to compute the adjoint lists of several simple examples.
@@ 1. Introduction
The foundation of our new approach to adjoints was initially developed in [9], where a family of independent sets of finite, atomic, and graded lattices was introduced and leveraged to prove an embedding theorem for lattices in which the family of independent sets constitutes a matroid. In this paper, we extend these results to cases where the family of independent sets does not constitute a matroid, allowing for further applications to the problem of adjoints.
The structure of this paper is as follows. Section 2 is dedicated to preliminaries on lattices ,matroids, and the adjoint matroid. In Section 3 we discuss the family of independent sets of a finite, atomic and graded lattice, recalling results from [9] and presenting a new embedding theorem. In Section 4 we show the effects of basic lattice operations on its family of independent sets and a class of special maximal independent sets. Section 5 moves to working with adjoint matroids centered around the characterization theorems of adjoints as an application of the family of independent sets. We further demonstrate how these methods can be applied to establish known results, specifically proving that every representable matroid and co-rank $3$ matroid possess an adjoint in a significantly simpler manner. Additionally, an immediate consequence is the identification of the explicit unique adjoint of a modular matroid. In Section 6 we present some partial results showing that the combinatorial derived matroid is an adjoint and discussing the existence of a maximal adjoint. Finally, this paper also includes an appendix in which we formulate an algorithm to check for adjoints based on the theorems in Section 5. We then use this algorithm to calculate the list of adjoints of some matroids of small rank and co-rank.
@@ 2. Preliminaries
We start with some basic lattice properties; for a more extensive background on matroids and geometric lattices, see [1]. A partially ordered set $\mathcal{L}$ in which every pair of elements has a unique supremum (also called join and denoted $\vee$) and a unique infimum (also called meet and denoted $\wedge$) is called a lattice. Two important elements, which always exist in a finite lattice, are the greatest element (denoted $\hat{1}$) and the least element (denoted $\hat{0}$). Given two elements $A\leq B$ in a lattice $\mathcal{L}$, we define the segment $A,B]_{\mathcal{L}}=\left\{C\mid A\leq C\leq B\right\}$. Notably, a segment of a lattice forms a lattice itself.
@@ 2.1. Definition:
We say $B$ covers $A$ in $\mathcal{L}$, and write $A\lessdot B$ if $A<B$ and for all $C$ such that $A\leq C\leq B$ we have $C=B$ or $C=A$. AA is an atom if it covers $\hat{0}$ and a coatom if it is covered by $\hat{1}$. A finite lattice is called atomic if every element is a join of atoms, and coatomic if its dual lattice is atomic. We denote the set of atoms of $\mathcal{L}$ by $A(\mathcal{L})$.
Finally, we say that a finite lattice is graded if there exists a rank function $\rho:\mathcal{L}\rightarrow\mathbb{N}$ with the following properties:
@@
    Compatibility with the order of $\mathcal{L}$, $A\leq B\iff\rho(A)\leq\rho(B)$.
@@
    Compatibility with the covering relation of $\mathcal{L}$, $A\lessdot B\Rightarrow\rho(A)=\rho(B)+1$.
@@ 2.2. Definition:
A geometric lattice is a finite graded lattice equipped with a submodular rank function, meaning it satisfies the inequality $\rho(A)+\rho(B)\leq\rho(A\vee B)+\rho(A\land B)$.
Next, we turn to some fundamental properties of matroids; for a more detailed discussion, refer to [1].
We adopt standard conventions: a matroid $\mathcal{M}$ refers to its set of bases. We denote the family of independent sets by $I(\mathcal{M})$, the collection of circuits by $\mathcal{C}(\mathcal{M})$, the associated geometric lattice by $\mathcal{L}(\mathcal{M})$, and the ground set by $E(\mathcal{M})$. When the reference to a specific matroid is clear, we omit $\mathcal{M}$, using, for instance, $\mathcal{L}$ instead of $\mathcal{L}(\mathcal{M})$.
An element $a\in E(\mathcal{M})$ will be called a loop of $\mathcal{M}$ if $\left\{a\right\}\notin I(\mathcal{M})$. Two elements $a,b\in E$ will be called parallel if both are not loops and $\left\{a,b\right\}\notin I(\mathcal{M})$ . We will say that a matroid is simple if it has no loops or parallel elements.
@@ 2.3. Definition:
Given two matroids $\mathcal{M}_{1},\mathcal{M}_{2}$ on the ground set $E$, we say $\mathcal{M}_{1}\leq\mathcal{M}_{2}$ in the weak order of matroids if $I(\mathcal{M}_{1})\subseteq I(\mathcal{M}_{2})$.
The dual matroid of $\mathcal{M}$, denoted $\mathcal{M}^{*}$ is the matroid defined on the same ground set, where its bases (maximal independent sets) are precisely the complements of the bases of $\mathcal{M}$. The geometric lattice associated with $\mathcal{M}^{*}$ is denoted by $\mathcal{L}^{*}$ and a loop of $\mathcal{M}^{*}$ is referred to as a co-loop.
A known property is that the complements of the hyperplanes (another name for co-atoms of geometric lattices) are precisely the circuits of the dual matroid. This property enables us to interpret an extension of the dual lattice of a geometric lattice, as discussed in the following sections, as a matroid with ground set $\mathcal{C}(\mathcal{M})$. We structure our definitions accordingly to reflect this perspective:
@@ 2.4. Definition:
Let $\mathcal{L},\mathcal{L}^{\triangle}$ be geometric lattices. We say $\mathcal{L}^{\triangle}$ is an adjoint of $\mathcal{L}^{*}$ if there exists a rank preserving order embedding $f:\mathcal{L}^{opp}\rightarrow\mathcal{L}^{\triangle}$ which restricts to a bijection from $A(\mathcal{L}^{opp})$ to $A(\mathcal{L}^{\triangle})$. For matroids $\mathcal{M},\mathcal{M}^{\triangle}$, we say that $\mathcal{M}^{\triangle}$ is an adjoint of $\mathcal{M}$ if $\mathcal{L}(\mathcal{M}^{\triangle})$ is an adjoint of $\mathcal{L}(\mathcal{M}^{*})$.
The following two properties are known for adjoints (for example, in [5]):
@@ 3.1. Definition:
Let $(\mathcal{L},\omega)$ be a pair of a finite lattice $\mathcal{L}$ and a partial order $\omega$ on its set of atoms $A(\mathcal{L})$. A nonempty set $D\subseteq A(\mathcal{L})$ of atoms is bounded below or $BB$ if there exist $a\in A(\mathcal{L})$ such that $a$ is a strict lower bound for all $d\in D$ in the order $\omega$ and $a\leq\vee D$ in $\mathcal{L}$.
@@
Given a finite lattice $\mathcal{L}$ and $P$ the set of all linear orders on $A(\mathcal{L})$ we define the family of independent sets $I(\mathcal{L})$ of $\mathcal{L}$ as follows:
[…]
We refer to the elements in $I(\mathcal{L})$ as independent sets of $\mathcal{L}$. Before continuing, we note the following immediate properties:
  1. (1)
    A subset of an independent set is itself independent. In other words, independent sets form an abstract simplicial complex.
  2. (2)
    A set containing a single atom is not $BB$ in any order.
  3. (3)
    sets containing at most two atoms are always independent. This is done by choosing any order where one of the two elements is first.
Example:
Let $\mathcal{L}\subseteq 2^{[4]}$ be the non-geometric lattice containing all subsets of size $0,1,4$ and $\left\{1,2\right\}$, ordered by inclusion. Following Definition 3.1, we see that $I(\mathcal{L})$ contains $\left\{\left\{1,2,3\right\},\left\{1,2,4\right\}\right\}$ and all subsets of size $\leq 2$.
The constructions in this section can be thought of as a generalization of the following observation, proven in [9].
@@ 3.2. Theorem:
If $\mathcal{L}$ is a geometric lattice, then $I(\mathcal{L})$ is the family of independent sets of the matroid corresponding to $\mathcal{L}$.
Next, we review key properties and definitions established in [9].
@@ 3.3. Definition:
Let $\mathcal{L}$ be a finite, atomic, and graded lattice, and $A\subseteq A(\mathcal{L})$. If $\vee A=X\in\mathcal{L}$ we say that $A$ spans $X$ and define $\rho(A)\coloneqq\rho(X)$. An independent set $A$ of $\mathcal{L}$ is called geometric if $\rho(A)=\left|A\right|$.
3.4. Lemma:
Let $\mathcal{L}$ be a finite, and graded lattice (not necessarily atomic), and $A\in I(\mathcal{L})$. There exists an atom $a\in A$ such that $\rho(A)\lneq\rho(A\setminus\left\{a\right\})$.
@@ 3.5. Theorem:
Let $\mathcal{L}$ be a finite, atomic, and graded lattice, and $A\in I(\mathcal{L})$. It is always true that $\left|A\right|\leq\rho(A)$. Moreover, if $\left|A\right|\neq\rho(A)$ then there exists an independent set $A\subsetneq B\in I(\mathcal{L})$ such that $\vee B=\vee A$ and $\left|B\right|=\rho(B)$. Consequently, the rank of an element equals the size of a maximal independent set contained in it.
The following embedding theorem was introduced in [9] and covers the case in which $I(\mathcal{L})$ is a family of independent sets of a matroid.
3.6. Theorem:
Let $\mathcal{L}$ be a finite, atomic and graded lattice. If $I(\mathcal{L})$ is a family of independent sets of a matroid, then there exists a rank-preserving order embedding $f:\mathcal{L}\rightarrow\mathcal{P}$ for which the restriction to the set of atoms $f\mid_{A(\mathcal{L})}:A(\mathcal{L})\rightarrow A(\mathcal{P})$ is a bijection. With $\mathcal{P}$ being the geometric lattice corresponding to $I(\mathcal{L})$.
We proceed with the formulation of an embedding theorem for cases where $I(\mathcal{L})$ does not constitute a family of independent sets of a matroid.
@@ 3.7. Lemma:
Let $\mathcal{L}$ be a finite and graded lattice (not necessarily atomic) and let $f^{\prime}:A(\mathcal{L})\rightarrow A(\mathcal{P})$ be a bijection between the sets of atoms of $\mathcal{L}$ and a geometric lattice $\mathcal{P}$. If $f^{\prime}$ can be extended to a rank-preserving order embedding $f:\mathcal{L}\rightarrow\mathcal{P}$, then $f(I(\mathcal{L}))\subseteq I(\mathcal{P})$.
@@
Let $f:\mathcal{L}\rightarrow\mathcal{P}$ be such an embedding and assume $f(I(\mathcal{L}))\nsubseteq I(\mathcal{P})$. We can choose a minimal $A\in f(I(\mathcal{L}))\setminus I(\mathcal{P})$. Let $a\in f^{-1}(A)$ with $\rho(f^{-1}(I)\setminus\left\{a\right\})\lneq\rho(f^{-1}(A))$ guaranteed by Lemma 3.4. As $A\setminus\left\{f(a)\right\}$ is an independent set of $\mathcal{P}$ and $A$ is not an independent set of $\mathcal{P}$, we must have
$f(a)\in\vee_{\mathcal{P}}(A\setminus\left\{f(a)\right\})\leq_{\mathcal{P}}f(\vee_{\mathcal{L}}(f^{-1}(A)\setminus\left\{a\right\}))$
, which is a contradiction to the choice of $a$.
@@ 3.8. Theorem:
Let $\mathcal{L}$ be a finite, atomic and graded lattice, and let $f^{\prime}:A(\mathcal{L})\rightarrow A(\mathcal{P})$ be a bijection between the set of atoms of $\mathcal{L}$ and the set of atoms of a geometric lattice $\mathcal{P}$ of the same rank. $f^{\prime}$ can be extended to a rank-preserving order embedding $f:\mathcal{L}\rightarrow\mathcal{P}$ if and only if $f(I(\mathcal{L}))\subseteq I(\mathcal{P})$ and $f(A)\notin I(\mathcal{P})$ for every $A\in 2^{A(\mathcal{L})}$ such that $\left|A\right|\lneq\rho\left(A\right)$.
@@
If there exists a function $f$ then Lemma 3.7 implies the first condition. The second condition holds as if $f(A)\in I(\mathcal{P})$ then $\rho(f(A))=\left|A\right|$ and so $f$ is rank-preserving.
In the other direction, we define $f(X)\coloneqq\vee_{\mathcal{P}}(f(A(X)))$. First, notice that $\rho(X)\leq\rho(f(X))$ for all $X\in\mathcal{L}$ as $f(X)$ contains the image of a geometric spanning set of $X$. To see $\rho(X)\geq\rho(f(X))$ for all $X\in\mathcal{L}$ observe that it cannot contain a larger independent set of $\mathcal{P}$ as it will contradict the second condition. To see $f$ is an embedding, let $X\in\mathcal{L}$, $a\in A(\mathcal{L})\setminus A(X)$ and $B$ be a geometric independent set spanning $X$. We have that $f(A)\cup\left\{f(a)\right\}\in I(\mathcal{P})$ is an independent set of size $\rho(X)+1$ in $f(X)$ which contradicts $f$ being rank-preserving.